OpenStudy (anonymous):

How do i solve f(x)=45-3x?

OpenStudy (anonymous):

u can differenciate it................

OpenStudy (anonymous):

What is one way?

OpenStudy (anonymous):

solve for what?

OpenStudy (anonymous):

I need to find the intercepts if i put it on a graph

OpenStudy (anonymous):

put x=0 and solve for y to get y-intercept and similar for x

OpenStudy (anonymous):

Ok, thanks

OpenStudy (anonymous):

let y=f(x) y=45-3x put y=0 to find x-intercept 0=45-3x 3x=45 x=15 now put x=0 to find t-intercept y=45-3(0) y=45

OpenStudy (anonymous):

beshak

OpenStudy (anonymous):

moon meri is right

OpenStudy (anonymous):

but she is using my idea

OpenStudy (anonymous):

Moon meri has a good answer, as does yours. She just solved the entire equation. But thanks for the help